# Nested subspace learning with flags

### v1

@@ A.1 An IRLS algorithm for robust subspace recovery
Iteratively reweighted least squares (IRLS) is a ubiquitous method to solve optimization problems involving $L^{p}$-norms. Motivated by the computation of the geometric median (Weiszfeld, 1937), IRLS is highly used to find robust maximum likelihood estimates of non-Gaussian probabilistic models (typically those containing outliers) and finds application in robust regression (Huber, 1964), sparse recovery (Daubechies et al., 2010) etc.
The recent fast median subspace (FMS) algorithm (Lerman and Maunu, 2018a), achieving state-of-the-art results in RSR uses an IRLS scheme to optimize the Least Absolute Deviation (LAD) (12). The idea is to first rewrite the LAD as
@@
$\mathcal{S}^{t+1}=\underset{\mathcal{S}\in\operatorname{Gr}(p,q)}{\operatorname{argmin}}\,\sum_{i=1}^{n}w_{i}(\mathcal{S}^{t})\left\lVert x_{i}-\Pi_{\mathcal{S}}x_{i}\right\rVert_{2}^{2}.$ (30)
This weighted least-squares problem enjoys a closed-form solution which relates to the eigenvalue decomposition of the weighted covariance matrix $\sum_{i=1}^{n}w_{i}(\mathcal{S}^{t})x_{i}{x_{i}}^{\top}$ (Vidal et al., 2016, Chapter 3.3).
We wish to derive an IRLS algorithm for the flag-tricked version of the LAD minimization problem (13). In order to stay close in mind to the recent work of Peng et al. (2023) who proved convergence of a general class of IRLS algorithms under some mild assumptions, we first rewrite (13) as
$\leavevmode\nobreak\ \underset{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}{\operatorname{argmin}}\,\sum_{i=1}^{n}\rho(r(\mathcal{S}_{1:d},x_{i})),$ (31)
where $r(\mathcal{S}_{1:d},x)=\left\lVert x-\Pi_{\mathcal{S}_{1:d}}x\right\rVert_{2}$ is the residual and $\rho(r)=|r|$ is the outlier-robust loss function. Following Peng et al. (2023), the IRLS scheme associated with (31) is:
$\begin{cases}w_{i}^{t+1}=\rho^{\prime}(r(\mathcal{S}_{1:d}^{t},x_{i}))/r(\mathcal{S}_{1:d}^{t},x_{i})=1/\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2},\\ (\mathcal{S}_{1:d})^{t+1}=\underset{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}{\operatorname{argmin}}\,\sum_{i=1}^{n}w_{i}^{t+1}\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2}^{2}.\end{cases}$ (32)
@@
$\underset{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}{\operatorname{argmin}}\,\sum_{i=1}^{n}w_{i}\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2}^{2}$ (33)
@@
Therefore, we are exactly in the same case as in Theorem 4, if we replace $XX^{\top}$ with the reweighted covariance matrix $\sum_{i=1}^{n}w_{i}x_{i}{x_{i}}^{\top}$. This does not change the result, so we conclude with the end of the proof of Theorem 4 (which itself relies on Szwagier and Pennec (2024)).
Hence, one gets an IRLS scheme for the LAD minimization problem. One can modify the robust loss function $\rho(r)=|r|$ by a Huber-like loss function to avoid weight explosion. Indeed, one can show that the weight $w_{i}:=1/\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2}$ goes to infinity when the first subspace $\mathcal{S}_{1}$ of the flag gets close to $x_{i}$ Therefore in practice, we take
@@
The final proposed scheme is given in Algorithm 3, named fast median flag (FMF), in reference to the fast median subspace algorithm of Lerman and Maunu (2018a).
@@
$X\in\mathbb{R}^{p\times n}$ (data), $\quad q_{1}<\dots<q_{d}$ (signature), $\quad t_{max}$ (max number of iterations), $\quad\eta$ (convergence threshold), $\quad\varepsilon$ (Huber-like saturation parameter)
@@
Remark 12
This is far beyond the scope of the paper, but we believe that the convergence result of Peng et al. (2023, Theorem 1) could be generalized to the FMF algorithm, due to the compactness of flag manifolds and the expression of the residual function $r$.
@@ A.2 Proof of Proposition 8
Let $\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})$ and $U_{1:d+1}:=[U_{1}|U_{2}|\dots|U_{d}|U_{d+1}]\in\mathcal{O}(p)$ be an orthogonal representative of $\mathcal{S}_{1:d}$. One has:
@@
$=\sqrt{\frac{1}{d^{2}}{x_{i}}^{\top}{\left(\sum_{k=1}^{d}(I_{p}-\Pi_{\mathcal{S}_{k}})\right)}^{2}x_{i}},$ (41)
$=\sqrt{\frac{1}{d^{2}}{x_{i}}^{\top}\left(\sum_{k=1}^{d+1}(k-1)U_{k}{U_{k}}^{\top}\right)^{2}x_{i}},$ (42)
@@
$=\sqrt{\sum_{k=1}^{d+1}\left(\frac{k-1}{d}\right)^{2}{x_{i}}^{\top}\left(U_{k}{U_{k}}^{\top}\right)x_{i}},$ (44)
$\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2}$ $=\sqrt{\sum_{k=1}^{d+1}\left(\frac{k-1}{d}\right)^{2}\left\lVert{U_{k}}^{\top}x_{i}\right\rVert_{2}^{2}},$ (45)
@@ B.1 A Newton method for multilevel trace ratio problems
We wish to solve the following optimization problem (cf. Equation 17):
$\underset{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}{\operatorname{argmax}}\,\frac{\operatorname{tr}\left(\Pi_{\mathcal{S}_{1:d}}A\right)}{\operatorname{tr}\left(\Pi_{\mathcal{S}_{1:d}}B\right)}.$ (46)
We draw from Ngo et al. (2012) (itself based on Wang et al. (2007); Jia et al. (2009)) to develop an efficient algorithm for solving this problem.
@@
Let $\phi\colon\operatorname{Fl}(p,q_{1:d})\ni\mathcal{S}_{1:d}\mapsto{\operatorname{tr}\left(\sum_{k=1}^{d}\Pi_{\mathcal{S}_{k}}A\right)}/{\operatorname{tr}\left(\sum_{k=1}^{d}\Pi_{\mathcal{S}_{k}}B\right)}$. One can see that $\phi$ is well defined on $\operatorname{Fl}(p,q_{1:d})$ if and only if $\operatorname{rank}(B)>p-q_{d}$, which we assume in the following. $\phi$ is continuous on $\operatorname{Fl}(p,q_{1:d})$ which is a smooth compact manifold, therefore it admits a maximum $\rho^{*}$. As a consequence, for any ${\mathcal{S}_{1:d}}\in\operatorname{Fl}(p,q_{1:d})$, one has $\phi(\mathcal{S}_{1:d})\leq\rho^{*}$, which is equivalent to $\leavevmode\nobreak\ {\sum_{k=1}^{d}\operatorname{tr}\left(\Pi_{\mathcal{S}_{k}}(A-\rho^{*}B)\right)\leq 0}$, and since the maximum is attained, lets say at $\mathcal{S}_{1:d}^{*}:=(\mathcal{S}_{1}^{*},\dots,\mathcal{S}_{d}^{*})$, one has ${\sum_{k=1}^{d}\operatorname{tr}\left(\Pi_{\mathcal{S}_{k}^{*}}(A-\rho^{*}B)\right)=0}$.
Using a similar reasoning as in the proof of Theorem 4 (itself based on Szwagier and Pennec (2024)), we can show that for any $\rho\in\mathbb{R}$, the function $\mathcal{S}_{1:d}\mapsto\sum_{k=1}^{d}\operatorname{tr}\left(\Pi_{\mathcal{S}_{k}}(A-\rho B)\right)$ is maximized in ${\mathcal{S}_{1:d}^{*}}=\left(\operatorname{Span}(v_{1},\dots,v_{q_{1}}),\operatorname{Span}(v_{1},\dots,v_{q_{2}}),\dots,\operatorname{Span}(v_{1},\dots,v_{q_{d}})\right)$, where $v_{1},\dots,v_{p}$ are the eigenvectors associated with the ordered-descending eigenvalues $\ell_{1},\dots,\ell_{p}$. The maximum is simply given by $f(\rho):=\sum_{k=1}^{d}\operatorname{tr}\left(\Pi_{\mathcal{S}_{k}^{*}}(A-\rho B)\right)=\sum_{k=1}^{d}\sum_{j=1}^{q_{k}}\ell_{j}$ and it verifies $f(\rho^{*})=0$.
@@
Let us now show that $\rho^{*}$ is the unique root of $f$. Let us introduce $\ell_{j}(\cdot)$, the function that maps a symmetric matrix to its $j$-th largest eigenvalue, and similarly $\mathcal{S}_{k}^{*}(\cdot)$, the function that maps a symmetric matrix to the subspace spanned by the eigenvectors associated with the $q_{k}$ largest eigenvalues. Then one can show that the derivative of $\sum_{j=1}^{q_{k}}\ell_{j}(A-\rho B)$ with respect to $\rho$ is $-\operatorname{tr}\left(\Pi_{\mathcal{S}_{k}^{*}(A-\rho B)}B\right)$ (Ngo et al., 2012). Therefore, the derivative of $f$ is $f^{\prime}(\rho)=-\sum_{k=1}^{d}\operatorname{tr}\left(\Pi_{\mathcal{S}_{k}^{*}(A-\rho B)}B\right)$, which is everywhere strictly negative since $B$ is positive semidefinite and $\operatorname{rank}(B)>p-q_{d}$. Therefore, f is a strictly decreasing function. Hence, $\rho^{*}$ is the unique root of $f$.
Now, the idea is to develop a Newton method to find the root $\rho^{*}$. The optimal flag will then be given by the eigenvalue decomposition of $A-\rho^{*}B$. One has the following iteration:
$\rho^{t+1}=\rho^{t}-\frac{f(\rho^{t})}{f^{\prime}(\rho^{t})}=\frac{\sum_{k=1}^{d}\operatorname{tr}\left(\Pi_{\mathcal{S}_{k}(A-\rho^{t}B)}A\right)}{\sum_{k=1}^{d}\operatorname{tr}\left(\Pi_{\mathcal{S}_{k}(A-\rho^{t}B)}B\right)}$ (47)
@@
Finally, in short, one can use similar considerations as in Ngo et al. (2012) to initialize the algorithm. If we assume that $B$ is positive definite, then we can find $Z\in\operatorname{GL}_{p}$ such that $Z^{\top}BZ=I_{p}$ and $Z^{\top}AZ=\Lambda$, where $\Lambda:=\operatorname{diag}\left(\lambda_{1}\geq\dots\geq\lambda_{p}\right)$ are the (ordered descending) generalized eigenvalues of the pencil $(A,B)$. Then, using Sylvester’s law of inertia, one gets that the number of negative eigenvalues of $Z^{\top}(A-\rho B)Z$ and of $\Lambda-\rho I_{p}$ are the same. Therefore, for $\rho=\lambda_{1}$, all the eigenvalues of $A-\rho B$ are nonpositive, so $f(\rho)=\sum_{k=1}^{d}\sum_{j=1}^{q_{k}}\lambda_{j}(A-\rho B)\leq 0$. Similarly, for $\rho=\lambda_{q_{d}}$, the $q_{d}$ largest eigenvalues of $A-\rho B$ are nonnegative, so in particular for all $k\in[1,d]$, the $q_{k}$ largest eigenvalues of $A-\rho B$ are nonnegative, so in the end $f(\rho)\geq 0$. To conclude, if $B$ is positive definite, then one has $\rho^{*}\in[\lambda_{q_{d}}(A,B),\lambda_{1}(A,B)]$. This can be used as an initialization to the Newton method, although in practice a random initialization is performed in practice in Ngo et al. (2012, Algorithm 4.1).
@@
The final proposed scheme is given in Algorithm 4, called flag iterative trace ratio in reference to the iterative trace ratio of Wang et al. (2007).
@@
$A,B\in\operatorname{SPD}_{p}$ (can find tighter restrictions), $\quad q_{1}<\dots<q_{d}$ (signature), $\quad\eta$ (convergence threshold)
@@
$t\leftarrow 0,\Delta\leftarrow\infty,\rho^{0}\in\left[\lambda_{q}(A,B),\lambda_{1}(A,B)\right]$
@@
     $U^{t},\Lambda^{t}\leftarrow\operatorname{EVD}(A-\rho^{t}B,q)$
     $\rho^{t}\leftarrow\frac{\sum_{k=1}^{d}\operatorname{tr}\left({U^{t}_{q_{k}}}^{\top}AU^{t}_{q_{k}}\right)}{\sum_{k=1}^{d}\operatorname{tr}\left({U^{t}_{q_{k}}}^{\top}BU^{t}_{q_{k}}\right)}$
     $\Delta\leftarrow\sqrt{\sum_{k=1}^{d}\Theta(U^{t}_{q_{k}},U^{t-1}_{q_{k}})^{2}}$

### v2

@@ A.1 An IRLS Algorithm for Robust Subspace Recovery
Iteratively reweighted least squares (IRLS) is a ubiquitous method to solve optimization problems involving $\ell^{p}$-norms. Motivated by the computation of the geometric median (weiszfeld_sur_1937), IRLS is highly used to find robust maximum likelihood estimates of non-Gaussian probabilistic models (typically those containing outliers) and finds application in robust regression (huber_robust_1964), sparse recovery (daubechies_iteratively_2010), etc.
The recent fast median subspace (FMS) algorithm (lerman_fast_2018), achieving state-of-the-art results in RSR uses an IRLS scheme to optimize the absolute deviation (12). The idea is to first rewrite the absolute deviation as
@@
$\mathcal{S}^{t+1}=\operatorname*{arg\,min}_{\mathcal{S}\in\operatorname{Gr}(p,q)}\sum_{i=1}^{n}w_{i}(\mathcal{S}^{t})\left\lVert x_{i}-\Pi_{\mathcal{S}}x_{i}\right\rVert_{2}^{2}.$ (28)
This weighted least-squares problem enjoys a closed-form solution which relates to the eigenvalue decomposition of the weighted covariance matrix $\sum_{i=1}^{n}w_{i}(\mathcal{S}^{t})x_{i}{x_{i}}^{\top}$ (vidal_generalized_2016, Chapter 3.3).
We wish to derive an IRLS algorithm for the flag-tricked version of the LAD problem (13). In order to stay close in mind to the recent work of peng_convergence_2023 who proved convergence of a general class of IRLS algorithms under some mild assumptions, we first rewrite (13) as
$\penalty 10000\ \operatorname*{arg\,min}_{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}\sum_{i=1}^{n}\rho(r(\mathcal{S}_{1:d},x_{i})),$ (29)
where $r(\mathcal{S}_{1:d},x)=\left\lVert x-\Pi_{\mathcal{S}_{1:d}}x\right\rVert_{2}$ is the residual and $\rho(r)=|r|$ is the outlier-robust loss function. Following peng_convergence_2023, the IRLS scheme associated with (29) is:
$\begin{cases}w_{i}^{t+1}=\rho^{\prime}(r(\mathcal{S}_{1:d}^{t},x_{i}))/r(\mathcal{S}_{1:d}^{t},x_{i})=1/\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2},\\ (\mathcal{S}_{1:d})^{t+1}=\operatorname*{arg\,min}_{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}\sum_{i=1}^{n}w_{i}^{t+1}\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2}^{2}.\end{cases}$ (30)
@@
$\operatorname*{arg\,min}_{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}\sum_{i=1}^{n}w_{i}\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2}^{2}$ (31)
@@
Therefore, we are exactly in the same case as in Theorem 8, if we replace $XX^{\top}$ with the reweighted covariance matrix $\sum_{i=1}^{n}w_{i}x_{i}{x_{i}}^{\top}$. This does not change the result, so we conclude with the end of the proof of Theorem 8 (itself relying on szwagier_curse_2025).
Hence, one gets an IRLS scheme for the LAD problem. One can modify the robust loss function $\rho(r)=|r|$ by a Huber-like loss function to avoid weight explosion. Indeed, one can show that the weight $w_{i}\coloneqq 1/\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2}$ goes to infinity when the first subspace $\mathcal{S}_{1}$ of the flag gets close to $x_{i}$ . Therefore in practice, we take
@@
The final proposed scheme is given in Algorithm 3, named fast median flag (FMF), in reference to the fast median subspace algorithm of lerman_fast_2018.
@@
$X\in\mathbb{R}^{p\times n}$ (data); $q_{1:d}\coloneqq(q_{1},\dots,q_{d})$ (signature); $t_{max}$ (max number of iterations); $\eta$ (convergence threshold); $\varepsilon$ (Huber-like saturation parameter)
@@
Remark 19 (Convergence Analysis)
This is far beyond the scope of the paper, but we believe that the convergence result of peng_convergence_2023 could be generalized to the FMF algorithm, due to the compactness of flag manifolds and the expression of the residual function $r$.
@@ A.2 Proof of Proposition 14
Let $\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})$ and $U_{1:d+1}\coloneqq\begin{bmatrix}U_{1}&U_{2}&\cdots&U_{d}&U_{d+1}\end{bmatrix}\in\mathcal{O}(p)$ be an orthogonal representative of $\mathcal{S}_{1:d}$ (cf. Section 2). One has:
@@
$=\sqrt{\frac{1}{d^{2}}{x_{i}}^{\top}{\left(\sum_{k=1}^{d}(I_{p}-\Pi_{\mathcal{S}_{k}})\right)}^{2}x_{i}}.$ (39)
By definition, one has $\Pi_{\mathcal{S}_{k}}=\sum_{l=1}^{k}U_{l}{U_{l}}^{\top}$ and, in particular, $I_{p}=\sum_{l=1}^{d+1}U_{l}{U_{l}}^{\top}$. Therefore, by re-injection and double summation inversion, one gets:
$\sum_{k=1}^{d}(I_{p}-\Pi_{\mathcal{S}_{k}})=\sum_{k=1}^{d}\left(\sum_{l=1}^{d+1}U_{l}U_{l}^{\top}-\sum_{l=1}^{k}U_{l}U_{l}^{\top}\right)=\sum_{k=1}^{d}\sum_{l=k+1}^{d+1}U_{l}U_{l}^{\top}=\sum_{l=2}^{d+1}\sum_{k=1}^{l-1}U_{l}U_{l}^{\top}=\\ \sum_{l=2}^{d+1}(l-1)\,U_{l}U_{l}^{\top}=\sum_{l=1}^{d+1}(l-1)\,U_{l}U_{l}^{\top}.$ (40)
Re-injecting the previous quantity in the computation of $\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2}$, one gets:
$\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2}$ $=\sqrt{\frac{1}{d^{2}}{x_{i}}^{\top}\left(\sum_{k=1}^{d+1}(k-1)U_{k}{U_{k}}^{\top}\right)^{2}x_{i}},$ (41)
@@
$=\sqrt{\sum_{k=1}^{d+1}\left(\frac{k-1}{d}\right)^{2}{x_{i}}^{\top}U_{k}{U_{k}}^{\top}x_{i}},$ (43)
$\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2}$ $=\sqrt{\sum_{k=1}^{d+1}\left(\frac{k-1}{d}\right)^{2}\left\lVert{U_{k}}^{\top}x_{i}\right\rVert_{2}^{2}}.$ (44)
Hence, the flag-tricked robust subspace recovery problem (13) is equivalent to the following problem on the orthogonal group:
$\operatorname*{arg\,min}_{U_{1:d+1}\in\mathcal{O}(p)}\sum_{i=1}^{n}\sqrt{\sum_{k=1}^{d+1}\left(\frac{k-1}{d}\right)^{2}\left\lVert{U_{k}}^{\top}x_{i}\right\rVert_{2}^{2}}.$ (45)
This problem can be further simplified into an optimization problem on Stiefel manifolds, by noting that:
$\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2}$ $=\sqrt{\sum_{k=1}^{d}\left(\left(\frac{k-1}{d}\right)^{2}\left\lVert{U_{k}}^{\top}x_{i}\right\rVert_{2}^{2}\right)+\left(\frac{d+1-1}{d}\right)^{2}\left\lVert{U_{d+1}}^{\top}x_{i}\right\rVert_{2}^{2}},$ (46)
$=\sqrt{\sum_{k=1}^{d}\left(\left(\frac{k-1}{d}\right)^{2}\left\lVert{U_{k}}^{\top}x_{i}\right\rVert_{2}^{2}\right)+{x_{i}}^{\top}{U_{d+1}}{U_{d+1}}^{\top}{x_{i}}},$ (47)
$=\sqrt{\sum_{k=1}^{d}\left(\left(\frac{k-1}{d}\right)^{2}\left\lVert{U_{k}}^{\top}x_{i}\right\rVert_{2}^{2}\right)+{x_{i}}^{\top}(I_{p}-{U_{1:d}}{U_{1:d}}^{\top}){x_{i}}},$ (48)
$=\sqrt{\sum_{k=1}^{d}\left(\left(\frac{k-1}{d}\right)^{2}\left\lVert{U_{k}}^{\top}x_{i}\right\rVert_{2}^{2}\right)+\left\lVert x_{i}\right\rVert_{2}^{2}-\left\lVert{U_{1:d}}^{\top}x_{i}\right\rVert_{2}^{2}},$ (49)
@@ Appendix B Trace Ratio Problems: Extensions and Proofs
In this section, we successively develop a Newton-Lanczos algorithm for the multilevel trace ratio problem (18), a kernelization of the same problem to handle nonlinear data, and we provide the proof of Proposition 16.
@@ B.1 A Newton Method for Multilevel Trace Ratio Problems
We wish to solve the following optimization problem (cf. (18)):
$\operatorname*{arg\,max}_{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}\frac{\operatorname{tr}\left(\Pi_{\mathcal{S}_{1:d}}A\right)}{\operatorname{tr}\left(\Pi_{\mathcal{S}_{1:d}}B\right)}.$ (50)
In this subsection, we propose a Newton method—drawn from ngo_trace_2012, which is itself based on wang_trace_2007 and jia_trace_2009—to solve this problem. The following proposition states that the multilevel trace ratio problem (50) can be reformulated as the search for the root (or zero) of a certain function, which can be then classically solved via Newton’s method.
Proposition 20 (Multilevel Trace Ratio as a Newton Method)
Let $M$ be a symmetric matrix, $(\ell_{j}(M),v_{j}(M))$ its $j$-th largest (eigenvalue, eigenvector)-pair and $\mathcal{S}_{k}^{*}(M)=\operatorname{Span}(v_{1}(M),\dots,v_{q_{k}}(M))$. Let $f(\rho)\coloneqq\sum_{k=1}^{d}\sum_{j=1}^{q_{k}}\ell_{j}(A-\rho B)$. Then $f$ admits a unique root $\rho^{*}$, and one has:
$\operatorname*{arg\,max}_{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}\frac{\operatorname{tr}\left(\Pi_{\mathcal{S}_{1:d}}A\right)}{\operatorname{tr}\left(\Pi_{\mathcal{S}_{1:d}}B\right)}=(\mathcal{S}_{1}^{*}(A-\rho^{*}B),\dots,\mathcal{S}_{d}^{*}(A-\rho^{*}B))$ (51)
@@
Proof Let $\phi\colon\operatorname{Fl}(p,q_{1:d})\ni\mathcal{S}_{1:d}\mapsto{\operatorname{tr}\left(\sum_{k=1}^{d}\Pi_{\mathcal{S}_{k}}A\right)}/{\operatorname{tr}\left(\sum_{k=1}^{d}\Pi_{\mathcal{S}_{k}}B\right)}$ be the objective function we want to maximize. We can see that $\phi$ is well defined on $\operatorname{Fl}(p,q_{1:d})$ if and only if $\operatorname{rank}(B)>p-q_{d}$, which we assume in the following.
$\phi$ is continuous on $\operatorname{Fl}(p,q_{1:d})$ which is a smooth compact manifold, therefore it admits a maximum $\rho^{*}$. As a consequence, for any ${\mathcal{S}_{1:d}}\in\operatorname{Fl}(p,q_{1:d})$, one has $\phi(\mathcal{S}_{1:d})\leq\rho^{*}$, which is equivalent to $\penalty 10000\ {\psi(\mathcal{S}_{1:d})\coloneqq\sum_{k=1}^{d}\operatorname{tr}\left(\Pi_{\mathcal{S}_{k}}(A-\rho^{*}B)\right)\leq 0}$. Similarly to Theorem 8, one can show that the maximum of $\psi$, which is 0, is attained in $\mathcal{S}_{1:d}^{*}\coloneqq\mathcal{S}_{1}^{*}(A-\rho^{*}B),\dots,\mathcal{S}_{d}^{*}(A-\rho^{*}B)$. Therefore, one has $\psi(\mathcal{S}_{1:d}^{*})=\sum_{k=1}^{d}\operatorname{tr}(\Pi_{\mathcal{S}_{k}^{*}(A-\rho^{*}B)}(A-\rho^{*}B))=f(\rho^{*})=0$.
@@
Let us now show that $\rho^{*}$ is the unique root of $f$. One can show that the derivative of $\sum_{j=1}^{q_{k}}\ell_{j}(A-\rho B)$ with respect to $\rho$ is $\operatorname{tr}(\Pi_{\mathcal{S}_{k}^{*}(A-\rho B)}B)$ (ngo_trace_2012). Therefore, the derivative of $f$ is $f^{\prime}(\rho)=-\sum_{k=1}^{d}\operatorname{tr}(\Pi_{\mathcal{S}_{k}^{*}(A-\rho B)}B)$, which is everywhere (strictly) negative since $B$ is positive semidefinite and $\operatorname{rank}(B)>p-q_{d}$. Therefore, f is a (strictly) decreasing function. Hence, $\rho^{*}$ is the unique root of $f$.
Newton’s method to find the root $\rho^{*}$ of $f$ is described by the following iteration:
$\rho^{t+1}=\rho^{t}-\frac{f(\rho^{t})}{f^{\prime}(\rho^{t})}=\frac{\sum_{k=1}^{d}\operatorname{tr}\left(\Pi_{\mathcal{S}_{k}(A-\rho^{t}B)}A\right)}{\sum_{k=1}^{d}\operatorname{tr}\left(\Pi_{\mathcal{S}_{k}(A-\rho^{t}B)}B\right)}.$ (52)
The optimal flag $\mathcal{S}_{1:d}^{*}$ is given by the eigenvalue decomposition of $A-\rho^{\infty}B$, according to Proposition 20. The only remaining question is the one of initialization. The following proposition provides a possible answer.
Proposition 21 (Bounds on Newton’s Root)
Let $\ell_{j}(A,B)$ denote the $j$-th largest eigenvalue of the matrix pencil $(A,B)$. If $B\succ 0$, then $\rho^{*}\in[\ell_{q_{d}}(A,B),\ell_{1}(A,B)]$.
@@
Proof Let $Z\in\operatorname{GL}_{p}$ such that $Z^{\top}BZ=I_{p}$ and $Z^{\top}AZ=\operatorname{diag}\left(\ell_{1}(A,B),\dots,\ell_{p}(A,B)\right)$. Then, using Sylvester’s law of inertia, one gets that the number of negative eigenvalues of $Z^{\top}(A-\rho B)Z$ and of $\Lambda-\rho I_{p}$ are the same. Therefore, for $\rho=\ell_{1}(A,B)$, all the eigenvalues of $A-\rho B$ are nonpositive, so $f(\rho)=\sum_{k=1}^{d}\sum_{j=1}^{q_{k}}\ell_{j}(A-\rho B)\leq 0$. Similarly, for $\rho=\ell_{q_{d}}(A,B)$, the $q_{d}$ largest eigenvalues of $A-\rho B$ are nonnegative, so in particular for all $k\in\{1,\dots,d\}$, the $q_{k}$ largest eigenvalues of $A-\rho B$ are nonnegative, so in the end $f(\rho)\geq 0$.
@@
The bounds given by Proposition 21 can be used to initialize Newton’s method. Alternatively, one can initialize the algorithm with a random flag (see Remark 4), similarly to ngo_trace_2012, in order to circumvent the potentially costly computation of the generalized eigenvalues. The final proposed scheme is given in Algorithm 4, called flag iterative trace ratio in reference to the iterative trace ratio of wang_trace_2007.
@@
$A,B\in\operatorname{Sym}_{+}(p)$; $q_{1:d}\coloneqq(q_{1},\dots,q_{d})$ (signature); $\eta$ (convergence threshold)
@@
$t\leftarrow 0,\Delta\leftarrow\infty,\rho^{0}\in\left[\ell_{q}(A,B),\ell_{1}(A,B)\right]$
@@
  $\begin{bmatrix}u_{1}^{t}&\cdots&u_{q}^{t}\end{bmatrix}\leftarrow\operatorname*{arg\,max}_{U\in\operatorname{St}(p,q)}U^{\top}(A-\rho^{t}B)U$ $\triangleright$ compute leading $q$ eigenvectors
  $\rho^{t}\leftarrow\frac{\sum_{k=1}^{d}\operatorname{tr}\left(\begin{bmatrix}u_{1}^{t}&\cdots&u_{q_{k}}^{t}\end{bmatrix}^{\top}A\begin{bmatrix}u_{1}^{t}&\cdots&u_{q_{k}}^{t}\end{bmatrix}\right)}{\sum_{k=1}^{d}\operatorname{tr}\left(\begin{bmatrix}u_{1}^{t}&\cdots&u_{q_{k}}^{t}\end{bmatrix}^{\top}B\begin{bmatrix}u_{1}^{t}&\cdots&u_{q_{k}}^{t}\end{bmatrix}\right)} \triangleright$ update trace ratio
  $\Delta\leftarrow\sqrt{\sum_{k=1}^{d}\Theta(\begin{bmatrix}u_{1}^{t}&\cdots&u_{q_{k}}^{t}\end{bmatrix},\begin{bmatrix}u_{1}^{t-1}&\cdots&u_{q_{k}}^{t-1}\end{bmatrix})^{2}}$ $\triangleright$ evaluate convergence